# Does Positive Reinforcement Work?: A Quasi-Experimental Study of the Effects of Positive Feedback on Reddit

### v1

@@ 6.5. Limitations and Future Work
Our research opens many avenues for the exploration of positive reinforcement and quantifying desirable content in online spaces. We summarize the limitations of our work and propose extensions to address these limitations.
@@
Our analysis does not consider what type of content is receiving the positive feedback, nor whether it is deserving of such approval. Desirability varies greatly across communities, thus further analysis is needed to model desirability within and across online communities.
6.5.2. Evaluate effectiveness of content-specific reinforcement.
Our analysis showed that recipients of positive feedback post content afterwards that receives even more positive feedback than they would have without the treatment. However, we do not know whether the future contributions are similar to the treated posts in content. Additionally, we limit our analysis to posts which typically have less text to analyze than comments. This research could be extended by exploring whether positive feedback reinforces content-level features of Reddit posts and comments.
@@ 6.5.4. Compare community feedback to moderator-specific forms of feedback
Our choice of treatments is limited in that some moderators do not consider upvotes a signal of quality (Jiang et al., 2023). This can be problematic in communities that require expertise for participation, since users giving out upvotes may not be equipped to evaluate how much a post deserves positive feedback. In such communities, moderators may be better judges of quality. Prior work shows that moderators are utilizing moderator-specific signals as a means for positive reinforcement (Lambert et al., 2024), thus future work should study how these forms of feedback affect the recipients and whether positive feedback from a position of authority has more or less capacity to reinforce user-level outcomes than anonymous forms of feedback.
@@ 7. Conclusion
In this work, we examined the effects of two forms of positive feedback on Reddit users and study whether positive reinforcement can be employed through distributed moderation practices. We found that both receiving a gild on a post and being highly-upvoted by the community encourages users to post content in the future that receives even more positive feedback than they would have otherwise. Through a user-specific analysis, we also find that receiving larger amounts of positive feedback is associated with larger effects on future positive feedback though users already satiated by recent positive feedback will not see as strong effects. Importantly, we also found that newcomers experience stronger effects of being highly-upvoted with respect to their future rate of contribution and norm acquisition. Based on our findings, we contribute design recommendations and discuss the implications of our work with respect to moderators, communities, and platforms beyond Reddit.

### v2

@@ 6.5. Limitations and Future Work
Our research opens many avenues for the exploration of positive reinforcement and quantifying desirable content in online spaces. We summarize the limitations of our work and propose extensions to address them.
@@
Our analysis does not consider whether the posts that received the positive feedback are deserving of such approval. Definitions of desirability can vary greatly (Lambert et al., 2024b; Goyal et al., 2024), thus future work should explore what is considered desirable within and across online communities, and develop context-sensitive approaches to model and identify desirable behavior online.
6.5.2. Extend understanding of content-specific reinforcement.
Our analysis showed that recipients of positive feedback receive even more positive feedback on subsequent posts, but that the content of those posts is more distinct from the rewarded post than it would have been without the treatment. This suggests that receiving positive feedback helps users apply community norms beyond making redundant contributions. However, this effect was small and it is possible that working with posts, which do not always have much text, does not allow us to capture content-level reinforcement. We suggest researchers extend this work by exploring the effects of positive feedback on reinforcing specific types of content within both posts and comments.
@@ 6.5.4. Compare community feedback to moderator-specific forms of feedback
Our choice of treatments is limited in that some moderators do not consider upvotes a signal of quality (Jiang et al., 2023). This can be problematic in communities that require expertise for participation, since users giving out upvotes may not be equipped to evaluate how much a post deserves positive feedback. In such communities, moderators may be better judges of quality. Prior work shows that moderators are utilizing moderator-specific signals as a means for positive reinforcement (Lambert et al., 2024a), thus future work should study how these forms of feedback affect the recipients and whether positive feedback from a position of authority has more or less capacity to reinforce user-level outcomes than anonymous forms of feedback.
@@ 7. Conclusion
In this work, we examined the effects of two forms of positive feedback on Reddit users and study whether positive reinforcement can be employed through distributed moderation practices. We found that both receiving gold on a post and being highly-upvoted by the community encourage users to post content in the future that receives even more positive feedback than they would have otherwise. Through a user-specific analysis, we also find that receiving larger amounts of positive feedback is associated with larger effects on future positive feedback though users already satiated by recent positive feedback will not see as strong effects. Importantly, we also found that newcomers experience stronger effects of being highly-upvoted with respect to their future rate of contribution and norm acquisition. Based on our findings, we contribute design recommendations and discuss the implications of our work with respect to moderators, communities, and platforms beyond Reddit.
Acknowledgements.
We thank the members of the Social Computing Lab (SCUBA) at the University of Illinois Urbana-Champaign for their feedback and input on this work. We also thank the anonymous reviewers for their thoughtful and valuable reviews.